Denote by a, b and c the column vectors a = (1 2 3), b = (-2 1 -3), c = (-2 -1 1) Calculate 2a - 5b, 2a- 5b +c, a'.b,
1
Expert's answer
2020-12-04T12:52:08-0500
"a=\begin{pmatrix}
 1 \\
 2 \\
3
\end{pmatrix}, b=\begin{pmatrix}
 -2 \\
 1 \\
-3
\end{pmatrix}, c=\begin{pmatrix}
 -2 \\
 -1 \\
1
\end{pmatrix}"

"2a-5b=2\begin{pmatrix}
 1 \\
 2 \\
3
\end{pmatrix}-5\begin{pmatrix}
 -2 \\
 1 \\
-3
\end{pmatrix}=\begin{pmatrix}
 2(1)-5(-2) \\
 2(2)-5(1) \\
2(3)-5(-3)
\end{pmatrix}"

"=\begin{pmatrix}
 12 \\
 -1 \\
21
\end{pmatrix}"

"2a-5b+c=2\begin{pmatrix}
 1 \\
 2 \\
3
\end{pmatrix}-5\begin{pmatrix}
 -2 \\
 1 \\
-3
\end{pmatrix}+\begin{pmatrix}
 -2 \\
 -1 \\
1
\end{pmatrix}"

"=\begin{pmatrix}
 12 \\
 -1 \\
21
\end{pmatrix}+\begin{pmatrix}
 -2 \\
 -1 \\
1
\end{pmatrix}=\begin{pmatrix}
 12-2 \\
 -1 -1 \\
21+1
\end{pmatrix}=\begin{pmatrix}
 10 \\
 -2 \\
22
\end{pmatrix}"

"a^T\cdot b=\begin{pmatrix}
 1 & 2 & 3
\end{pmatrix}\cdot\begin{pmatrix}
 -2 \\
 1 \\
-3
\end{pmatrix}"

"=1(-2)+2(1)+3(-3)=-9"